# Time-Series Foundation Model for Value-at-Risk

### v3

@@
Time-Series Foundation Model for Value-at-Risk
[…]
Time-Series Foundation Model for Value-at-Risk
[…]
This study is the first to explore the performance of a time-series foundation model for Value-at-Risk (VaR) estimation. Foundation models, pre-trained on vast and varied datasets, can be used in a zero-shot setting with relatively minimal data or further improved through finetuning. We compare the performance of Google’s model, called TimesFM, against conventional parametric and non-parametric models, including GARCH, Generalized Autoregressive Score (GAS), and empirical quantile estimates, using daily returns from the S&P 100 index and its constituents over 19 years. Our backtesting results indicate that in terms of the actual-over-expected ratio, the fine-tuned TimesFM model consistently outperforms traditional methods. Regarding the quantile score loss function, it achieves performance comparable to the best econometric approach, the GAS model. Overall, the foundation model is either the best or among the top performers in forecasting VaR across the 0.01, 0.025, 0.05, and 0.1 VaR levels. Fine-tuning significantly improves accuracy, indicating that zero-shot use is not optimal for VaR estimation.
[…]
Value at Risk, AI, Foundation models, TimesFM, Time series forecasting
@@ 1 Introduction
While Machine Learning (ML) has significantly advanced financial market research—such as in asset pricing, volatility forecasting, limit order book dynamics, and optimal hedging (Gu et al. 2020, Christensen et al. 2023, Tran et al. 2018, Mikkilä and Kanniainen 2023)—the risk management literature still largely relies on traditional econometric techniques, especially in the area of Value-at-Risk (VaR). One reason for this may be that Deep Learning (DL) models require substantial amounts of data, which poses a significant limitation when applying them to calculate risk metrics for individual stocks. VaR estimations are often based on daily return observations (for example, see González-Rivera et al. 2004, Patton et al. 2019, Taylor 2019, based solely on historical daily returns), and even 20 years of data would yield only around 5,000 observations. This is quite limited, especially considering that DL models often contain millions of parameters that need to be learned.
@@
However, a solution to this problem is currently underway. Recently, ML has experienced a paradigm shift with the emergence of foundation models, which are large-scale, general-purpose pre-trained and often fine-tunable models on diverse datasets spanning various data distributions. Although foundation models are based on deep learning and transfer learning principles, their scale leads to the emergence of new capabilities (Bommasani et al. 2021). In 2023–2024, following major advancements in large language models (LLMs), several pre-trained time-series foundation models have emerged, including Google’s TimesFM (Das et al. 2024), Time-GPT (Garza and Mergenthaler-Canseco 2024), and LagLlama (Rasul et al. 2023).
@@
AI-driven advancements in financial risk modeling have opened doors to new methodologies for assessing complex financial data (Tran et al. 2019, Guijarro-Ordonez et al. 2021, Chen et al. 2024). Historically, methods like GARCH-based models have been effective for capturing dependencies and volatility clustering in market data (Hoga and Demetrescu 2023). Recent approaches, however, such as Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), and Conditional GANs (CGANs), enhance distribution forecasting by generating realistic synthetic data (Nareklishvili et al. 2023, Athey et al. 2021, Hofert et al. 2022). Such models provide flexible, non-parametric approaches for data structures that traditional econometric tools may struggle to approximate. Ericson et al. (2024) presents a comparative review showing that deep generative methods outperform traditional benchmarks in modeling complex financial dynamics for VaR estimation. These advancements highlight the potential of AI-driven approaches in risk forecasting. This study stands out from previous research by adopting a time-series foundation model, pre-trained on diverse datasets, to forecast VaR. Unlike traditional supervised ML models trained for specific tasks, time-series foundation models are pre-trained on vast datasets spanning billions of time points. A single pre-trained model can be applied across a wide range of domains, from traffic forecasting to stock price predictions, offering quick and cost-effective solutions. The term ‘foundational’ emphasizes their flexibility, as they can be adapted for different tasks or domains with minimal additional training. The foundation models can be used with or without fine-tuning.
@@
We implement a pre-trained time-series foundation model, specifically Google’s TimesFM, for VaR estimation and compare it with existing state-of-the-art parametric and semi-parametric approaches, including GARCH model, Generalized Autoregressive Score (GAS) one-factor model, and empirical quantile estimates. The key question is whether the pre-trained foundation model yields results comparable to, or even capable of outperforming, existing econometric methods. An additional objective is to assess the extent to which fine-tuning improves the accuracy of VaR predictions. While several foundation models have been recently introduced, this study focuses on TimesFM due to its open-source availability [^1] https://github.com/google-research/TimesFM., modifiability, and free accessibility, ensuring the results are fully replicable. [^2] The codes of this paper can be found at https://github.com/Anubha0812/TimesFM-for-Value-at-Risk. Furthermore, the authors of TimesFM have empirically demonstrated that their model outperforms TimeGPT in forecasting accuracy, adding further motivation for its selection. Rather than identifying the best foundation model, which is subject to frequent updates and new developments, this study aims to address the fundamental question of whether state-of-the-art parametric and semi-parametric methods can be outperformed.
@@
Our main finding is that, in terms of the actual-over-expected ratio, the fine-tuned time-series foundation model outperforms existing econometric approaches for VaR estimation. Specifically, the fine-tuned foundation model consistently demonstrates superior performance, on average, across all 91 constituents of the S&P 100 included in the analysis. Regarding the quantile score loss function, which is the second backtesting metric, the fine-tuned foundation model achieves performance comparable to the best econometric approach, the GAS model. Overall, the foundation model is either the top performer or among the best in forecasting VaR across the 0.01, 0.025, 0.05, and 0.1 VaR levels. Moreover, we found that fine-tuning is practically necessary for the foundation model; using it as pre-trained in zero-shot settings results in significantly worse performance compared to the best econometric approaches. We consider this as a remarkable result on the performance of Google’s TimesFM model as this model is not explicitly designed for this purpose and was not even pre-trained with stock return data. This problem is statistically challenging due to its focus on heteroscedastic processes with time-varying conditional volatility. In this light, the paper is valuable not only for financial risk management, but also as a benchmark for testing AI time-series models against state-of-the-art statistical approaches on a highly complex issue.
The paper is structured as follows: In Section 2, we describe the time-series foundation model, Google’s TimesFM, and the benchmark models against which it is compared. Section 3 details the data and experimental setup for the models and fine-tuning specifications for VaR estimation. The evaluation metrics and backtesting procedures are outlined in Section 4. Section 5 presents an analysis of the empirical results, examining the performance of the foundation model relative to traditional econometric approaches. Finally, the findings have several implications beyond Value-at-Risk estimation techniques, which, along with the limitations of our analysis, will be discussed at the end of the paper in Section 6.
@@ 2.1 Time-Series Foundation Models
TimesFM model (Das et al. 2024) is a decoder-only transformer architecture, similar to GPT models. In contrast to encoder-decoder models, where the encoder processes the input sequence and the decoder generates output based on the encoded input, decoder-only models focus solely on output generation processing the input through successive decoder layers. This allows for faster training and inference times compared to encoder-decoder models. Decoder-only models like are highly advantageous for zero-shot forecasting, where minimal fine-tuning is required. On the other hand, the advantage of the encoder-decoder architecture, which is used, for example, in TimeGPT (Garza and Mergenthaler-Canseco 2024), is that it can handle more complex inputs and exogenous variables. However, in this paper, we use univariate and well-structured time-series data on returns, and VaR is calculated separately for each security, and for that reason, the encoder component does not add significant value.
@@
In pretraining, TimesFM was trained on a vast corpus comprising both real-world time-series data such as Google trends, Wiki Pageviews, Synthetic Data, and Electricity, Traffic, and Weather datasets used in M4 competition (Makridakis et al. 2020, Spiliotis et al. 2020). However, financial data was not used, and this aspect makes the application of the present paper particularly interesting. This corpus, although smaller in scale compared to text-based models, spans around 100 billion time points, providing the diversity needed for the model to generalize well across domains and forecasting tasks. Compared to the state-of-the-art LLMs, TimesFM time-series foundation model is much smaller, with 200M parameters and pretraining data consisting of around 100B timepoints. Despite its smaller scale, it was reported to achieve relatively high zero-shot forecasting accuracy.
@@ 2.2 Benchmark Models for VaR
    One of the most widely used approaches for VaR estimation is the rolling window method, which computes the risk measure based on historical data over a fixed time window:
@@
  • ARMA-GARCH Models (Angelidis et al. 2004):
@@
    3. 3.
      Empirical Distribution Function: As a nonparametric alternative, we use the Empirical Distribution Function (EDF) method (Manganelli and Engle 2004). This approach estimates the distribution of standardized residuals $\eta_{t}$ from historical data, avoiding any parametric assumptions.
  • One factor Gas model (Patton et al. 2019): This model belongs to the class of CAViaR models (Taylor 2005), where the evolution of the risk measure is directly parameterized without the need to specify a full conditional distribution for asset returns. [^3] GAS model was originally proposed in Creal et al. (2013). Patton et al. (2019) proposed a dynamic GAS model employing the class of loss functions proposed in Fissler and Ziegel (2016), called FZ loss functions. In this paper, we adopt the GAS model from Section 2.3 in Patton et al. (2019). The evolution of VaR for the Gas 1 factor model is given
@@ 3.1 Experiment Setup
We implement the recently introduced pre-trained time-series model, Google’s TimesFM, for forecasting Value at risk and compare its performance against state-of-the-art parametric and non-parametric approaches, namely, GARCH model with the standardized residual modeled as empirical distribution function (G-EDF), normal distribution (G-N) and Student’s-t (G-t) distribution, one factor Generalized AutoRegressive Score model (GAS) and rolling window method (Historical).
@@
For GARCH and GAS models, we followed Patton et al. (2019) and used January 2005 to December 2014 as a training period for estimating model parameters, which were then used to forecast over the test period from the 2nd January 2015 to the 19th September 2023, thus resulting in a sequence of 2,268 VaR estimates. For the rolling window (Historical) method, we used the data of 512 days preceding the first test day (i.e., 2nd January 2015), and forecasted the VaR at various levels. We then shift the training sample by one day forward in time to include more recent data and exclude the oldest data point such that a fixed size of the rolling window is maintained. At each rolling step, we obtain the prediction for the next day using the most recent 512 observations, thus resulting in a sequence of 2,268 VaR estimates for each level. For each of the models, G-EDF, G-N, G-t, one factor GAS, and Historical, we obtained the VaR estimates at levels 1%, 2.5%, 5%, and 10%.
@@
TimesFM is originally trained on a large number of public and synthetic datasets and exhibits robust out-of-the-box zero-shot performance in comparison to the accuracy of various state-of-the-art forecasting models specific to individual datasets under consideration. The model is flexible enough to forecast with different input length, prediction length and time granularities at inference time. The pre-trained model returns the point forecasts along with deciles, i.e., the quantiles [^5] As per Google’e TimesFM github page, “We experimentally offer quantile heads but they have not been calibrated after pretraining.” at $10\%,20\%,\ldots,90\%$ levels. We experimented with different prediction horizons with a fixed input length of 2 years (512 days). Specifically, we set the prediction horizon to 1 day, 21 days and 63 days and denote these models as PT1, PT21, and PT63. For PT1, the input window shifts daily with the prediction points, while for PT21 and PT63, we use a fixed set of input observations for daily forecasts over the subsequent 21 and 63 days, respectively. For example, with PT21 (or PT63), we estimated one-day VaRs over a 21-day (or 63-day) period starting from January 2, 2015, using the most recent 512 observations. After this, we shifted the training data forward by 21 days (or 63 days) to incorporate more recent data and applied the pre-trained model for the next 21 (or 63) days, generating one-day VaRs estimates over this period. The variation in prediction horizon allows us to examine the model’s ability to generalize across forecasting horizons of varying lengths, i.e., short and long horizons. However, as mentioned above, the pre-trained models, PT1, PT21, and PT63, only produced VaR forecast at 90% level.
@@ 4 Evaluation metrics
In financial risk management, VaR is a widely used measure that estimates the potential loss in the value of a stock over a given time period, with a certain confidence level. However, VaR estimates are erroneous and hence often biased due to various sources of errors such as model errors, sampling errors, etc. Therefore, to ensure that VaR models are reliable, it is crucial to perform rigorous backtesting. Backtesting is the process of comparing the VaR predictions against actual observed outcomes to determine the accuracy of the model. Since a VaR value is fundamentally unobservable, we need to judge the performance of a model by checking whether these estimates are consistent with the realized returns or not, given the confidence level at which these estimates were obtained. Thus, VaR backtesting is critical for evaluating the accuracy of the VaR forecasts.
@@
An alternative to AE ratio is the proportion of failure test (POF), also known as the unconditional coverage test (UC). which statistically evaluates whether the actual number of VaR violations matches the expected number of violations over time, based on the chosen confidence level. For instance, it calculates the failure rate defined as the proportion of number of violations ($A$) to the total number of observation $T$ and examines its consistency with the models confidence level $\alpha$. More specifically, let $I_{t}$ denotes the indicator function that the realized return was less than the VaR estimate, i.e.,
@@
As stated in Kupiec (1995), the UC test is capable of rejecting a model for both low and high failures, but has poor power in general. Further, the UC test tests the number of violations, while it is crucial to test whether these losses are spread evenly over time or not. The Conditional Coverage (CC) test, proposed by Christoffersen (1998), does the same, that is, it tests whether the violations are randomly distributed over time and not clustered. That is, CC test takes into account any conditionality in VaR estimates: the model must respond to clustering events such as if volatilities are high in some period and low in others. The test statistics is given by
@@
In the evaluation of VaR forecasts, the quantile score is a commonly used metric and is now standard in the VaR literature (Ehm et al. 2016, Gneiting 2011, Taylor 2019). The above mentioned statistical tests are good to evaluate the statistical adequacy of VaR estimates, that is, a model for VaR forecasting is deemed to be an adequate model if we fail to reject the null hypothesis. However, these tests cannot conclude whether an ‘adequate’ model outperforms another ‘adequate’ model. Scoring functions allow us to do an honest assessment of VaR as a perfect VaR estimate would always minimize the scoring function. The quantile score $QS_{t}$ can be defined as follows:
@@
For every instant $t$, a model’s score increases either by (a) $\alpha-1$ fraction of the distance between the realized return and the VaR forecast, if realized return is smaller than the VaR forecast for the same period, or (b) by $\alpha$ fraction of the distance otherwise. This scoring rule is consistent for quantile estimation, meaning that it achieves its minimum value when the quantile forecast $q_{t}$ correctly estimates the true quantile of the distribution. Therefore, models that produce lower average quantile scores in the out-of-sample period are considered superior. Finally, to compare the performance of two models, we use the Diebold–Mariano (DM) test (Diebold and Mariano 2002) with

### v4

@@
Abstract
This study is the first to explore the performance of a time-series foundation model for Value-at-Risk (VaR) forecasting. Foundation models, pre-trained on vast and varied datasets, can be used in a zero-shot setting with relatively minimal data or further improved through finetuning. We compare the performance of Google’s model, called TimesFM, against conventional parametric and non-parametric models, including GARCH, Generalized Autoregressive Score (GAS), and empirical quantile estimates, using daily returns from the S&P 100 index and its constituents over 19 years. Our backtesting results indicate that in terms of the actual-over-expected ratio, the fine-tuned TimesFM model consistently outperforms traditional methods. Regarding the quantile score loss function, it achieves performance comparable to the best econometric approach, the GAS model. Overall, the foundation model is either the best or among the top performers in forecasting VaR across the 0.01, 0.025, 0.05, and 0.1 VaR levels. Fine-tuning significantly improves accuracy, indicating that zero-shot use is not optimal for VaR forecasting.
keywords:
Econometric models , AI , Value-at-Risk Forecasting , TimesFM , Time series , Foundation Models
journal: International Journal of Forecasting
@@ Abstract
[inst1]organization=Tampere University, Department of Computing Sciences, Financial Computing and Data Analytics Group,addressline=Kalevantie 4, city=Tampere, postcode=33100, country=Finland
@@
[inst2]organization=IIT Ropar, Department of Mathematics,addressline=Ropar, city=Punjab, postcode=140001, country=India
@@ 1 Introduction
While Machine Learning (ML) has significantly advanced financial market research—such as in asset pricing, volatility forecasting, limit order book dynamics, and optimal hedging (Gu et al., 2020; Christensen et al., 2023; Tran et al., 2018; Mikkilä and Kanniainen, 2023)—the risk management literature still largely relies on traditional econometric techniques, especially in the area of Value-at-Risk (VaR). One reason for this may be that Deep Learning (DL) models require substantial amounts of data, which poses a significant limitation when applying them to calculate risk metrics for individual stocks. VaR forecasts are often based on daily return observations (for example, see González-Rivera et al., 2004; Patton et al., 2019; Taylor, 2019; Le, 2020; Taylor, 2020, based solely on historical daily returns), and even 20 years of data would yield only around 5,000 observations. This is quite limited, especially considering that DL models often contain millions of parameters that need to be learned.
@@
However, a solution to this problem is currently underway. Recently, ML has experienced a paradigm shift with the emergence of foundation models, which are large-scale, general-purpose pre-trained and often fine-tunable models on diverse datasets spanning various data distributions. Although foundation models are based on deep learning and transfer learning principles, their scale leads to the emergence of new capabilities (Bommasani et al., 2021). In 2023–2024, following major advancements in large language models (LLMs), several pre-trained time-series foundation models have emerged, including Google’s TimesFM (Das et al., 2024), Time-GPT (Garza and Mergenthaler-Canseco, 2024), and LagLlama (Rasul et al., 2023).
@@
AI-driven advancements in financial risk modeling have opened doors to new methodologies for assessing complex financial data (Tran et al., 2019; Guijarro-Ordonez et al., 2021; Chen et al., 2024). Historically, methods like GARCH-based models have been effective for capturing dependencies and volatility clustering in market data (Hoga and Demetrescu, 2023). Recent approaches, however, such as Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), and Conditional GANs (CGANs), enhance distribution forecasting by generating realistic synthetic data (Nareklishvili et al., 2023; Athey et al., 2021; Hofert et al., 2022). Such models provide flexible, non-parametric approaches for data structures that traditional econometric tools may struggle to approximate. Ericson et al. (2024) presents a comparative review showing that deep generative methods outperform traditional benchmarks in modeling complex financial dynamics for VaR forecasting. These advancements highlight the potential of AI-driven approaches in risk forecasting. This study stands out from previous research by adopting a time-series foundation model, pre-trained on diverse datasets, to forecast VaR. Unlike traditional supervised ML models trained for specific tasks, time-series foundation models are pre-trained on vast datasets spanning billions of time points. A single pre-trained model can be applied across a wide range of domains, from traffic forecasting to stock price predictions, offering quick and cost-effective solutions. The term ‘foundational’ emphasizes their flexibility, as they can be adapted for different tasks or domains with minimal additional training. The foundation models can be used with or without fine-tuning.
@@
We implement a pre-trained time-series foundation model, specifically Google’s TimesFM, for VaR forecasting and compare it with existing state-of-the-art parametric and semi-parametric approaches, including GARCH model, Generalized Autoregressive Score (GAS) one-factor model, and empirical quantile estimates. The key question is whether the pre-trained foundation model yields results comparable to, or even capable of outperforming, existing econometric methods. An additional objective is to assess the extent to which fine-tuning improves the accuracy of VaR predictions. While several foundation models have been recently introduced, this study focuses on TimesFM due to its open-source availability [^1] https://github.com/google-research/TimesFM., modifiability, and free accessibility, ensuring the results are fully replicable. [^2] The codes of this paper can be found at https://github.com/Anubha0812/TimesFM-for-Value-at-Risk. Furthermore, the authors of TimesFM have empirically demonstrated that their model outperforms TimeGPT in forecasting accuracy, adding further motivation for its selection. Rather than identifying the best foundation model, which is subject to frequent updates and new developments, this study aims to address the fundamental question of whether state-of-the-art parametric and semi-parametric methods can be outperformed.
@@
Our main finding is that, in terms of the actual-over-expected ratio, the fine-tuned time-series foundation model outperforms existing econometric approaches for VaR forecasting. Specifically, the fine-tuned foundation model consistently demonstrates superior performance, on average, across all 91 constituents of the S&P 100 included in the analysis. Regarding the quantile score loss function, which is the second backtesting metric, the fine-tuned foundation model achieves performance comparable to the best econometric approach, the GAS model. Overall, the foundation model is either the top performer or among the best in forecasting VaR across the 0.01, 0.025, 0.05, and 0.1 VaR levels. Moreover, we found that fine-tuning is practically necessary for the foundation model; using it as pre-trained in zero-shot settings results in significantly worse performance compared to the best econometric approaches. We consider this as a remarkable result on the performance of Google’s TimesFM model as this model is not explicitly designed for this purpose and was not even pre-trained with stock return data. This problem is statistically challenging due to its focus on heteroscedastic processes with time-varying conditional volatility. In this light, the paper is valuable not only for financial risk management, but also as a benchmark for testing AI time-series models against state-of-the-art statistical approaches on a highly complex issue.
The paper is structured as follows: In Section 2, we describe the time-series foundation model, Google’s TimesFM, and the benchmark models against which it is compared. Section 3 details the data and experimental setup for the models and fine-tuning specifications for VaR forecasting. The evaluation metrics and backtesting procedures are outlined in Section 4. Section 5 presents an analysis of the empirical results, examining the performance of the foundation model relative to traditional econometric approaches. Finally, the findings have several implications beyond Value-at-Risk forecasting techniques, which, along with the limitations of our analysis, will be discussed at the end of the paper in Section 6.
@@ 2.1 Time-Series Foundation Models
TimesFM model (Das et al., 2024) is a decoder-only transformer architecture, similar to GPT models. In contrast to encoder-decoder models, where the encoder processes the input sequence and the decoder generates output based on the encoded input, decoder-only models focus solely on output generation processing the input through successive decoder layers. This allows for faster training and inference times compared to encoder-decoder models. Decoder-only models like are highly advantageous for zero-shot forecasting, where minimal fine-tuning is required. On the other hand, the advantage of the encoder-decoder architecture, which is used, for example, in TimeGPT (Garza and Mergenthaler-Canseco, 2024), is that it can handle more complex inputs and exogenous variables. However, in this paper, we use univariate and well-structured time-series data on returns, and VaR is calculated separately for each security, and for that reason, the encoder component does not add significant value.
@@
In pretraining, TimesFM was trained on a vast corpus comprising both real-world time-series data such as Google trends, Wiki Pageviews, Synthetic Data, and Electricity, Traffic, and Weather datasets used in M4 competition (Makridakis et al., 2020; Spiliotis et al., 2020). However, financial data was not used, and this aspect makes the application of the present paper particularly interesting. This corpus, although smaller in scale compared to text-based models, spans around 100 billion time points, providing the diversity needed for the model to generalize well across domains and forecasting tasks. Compared to the state-of-the-art LLMs, TimesFM time-series foundation model is much smaller, with 200M parameters and pretraining data consisting of around 100B timepoints. Despite its smaller scale, it was reported to achieve relatively high zero-shot forecasting accuracy.
@@ 2.2 Benchmark Models for VaR
    One of the most widely used approaches for VaR forecasting is the rolling window method, which computes the risk measure based on historical data over a fixed time window:
@@
  • 2.
    ARMA-GARCH Models (Angelidis et al., 2004):
@@
    3. (c)
      Empirical Distribution Function: As a nonparametric alternative, we use the Empirical Distribution Function (EDF) method (Manganelli and Engle, 2004). This approach estimates the distribution of standardized residuals $\eta_{t}$ from historical data, avoiding any parametric assumptions.
  • 3.
    One factor Gas model (Patton et al., 2019): This model belongs to the class of CAViaR models (Taylor, 2005), where the evolution of the risk measure is directly parameterized without the need to specify a full conditional distribution for asset returns. [^3] GAS model was originally proposed in Creal et al. (2013). Patton et al. (2019) proposed a dynamic GAS model employing the class of loss functions proposed in Fissler and Ziegel (2016), called FZ loss functions. In this paper, we adopt the GAS model from Section 2.3 in Patton et al. (2019). The evolution of VaR for the Gas 1 factor model is given
@@ 3.1 Experiment Setup
We implement the recently introduced pre-trained time-series model, Google’s TimesFM, for forecasting Value-at-risk and compare its performance against state-of-the-art parametric and non-parametric approaches, namely, GARCH model with the standardized residual modeled as empirical distribution function (G-EDF), normal distribution (G-N) and Student’s-t (G-t) distribution, one factor Generalized AutoRegressive Score model (GAS) and rolling window method (Historical).
@@
For GARCH and GAS models, we followed Patton et al. (2019) and used January 2005 to December 2014 as a training period for estimating model parameters, which were then used to forecast over the test period from the 2nd January 2015 to the 19th September 2023, thus resulting in a sequence of 2,268 VaR forecasts. For the rolling window (Historical) method, we used the data of 512 days preceding the first test day (i.e., 2nd January 2015), and forecasted the VaR at various levels. We then shift the training sample by one day forward in time to include more recent data and exclude the oldest data point such that a fixed size of the rolling window is maintained. At each rolling step, we obtain the prediction for the next day using the most recent 512 observations, thus resulting in a sequence of 2,268 VaR forecasts for each level. For each of the models, G-EDF, G-N, G-t, one factor GAS, and Historical, we obtained the VaR forecasts at levels 1%, 2.5%, 5%, and 10%.
@@
TimesFM is originally trained on a large number of public and synthetic datasets and exhibits robust out-of-the-box zero-shot performance in comparison to the accuracy of various state-of-the-art forecasting models specific to individual datasets under consideration. The model is flexible enough to forecast with different input length, prediction length and time granularities at inference time. The pre-trained model returns the point forecasts along with deciles, i.e., the quantiles [^5] As per Google’e TimesFM GitHub page, “We experimentally offer quantile heads but they have not been calibrated after pretraining.” at $10\%,20\%,\ldots,90\%$ levels. We experimented with different prediction horizons with a fixed input length of 2 years (512 days). Specifically, we set the prediction horizon to 1 day, 21 days and 63 days and denote these models as PT1, PT21, and PT63. For PT1, the input window shifts daily with the prediction points, while for PT21 and PT63, we use a fixed set of input observations for daily forecasts over the subsequent 21 and 63 days, respectively. For example, with PT21 (or PT63), we estimated one-day VaRs over a 21-day (or 63-day) period starting from January 2, 2015, using the most recent 512 observations. After this, we shifted the training data forward by 21 days (or 63 days) to incorporate more recent data and applied the pre-trained model for the next 21 (or 63) days, generating one-day VaRs forecasts over this period. The variation in prediction horizon allows us to examine the model’s ability to generalize across forecasting horizons of varying lengths, i.e., short and long horizons. However, as mentioned above, the pre-trained models, PT1, PT21, and PT63, only produced VaR forecast at 90% level.
@@ 4 Evaluation metrics
In financial risk management, VaR is a widely used measure that estimates the potential loss in the value of a stock over a given time period, with a certain confidence level. However, VaR forecasts are erroneous and hence often biased due to various sources of errors such as model errors, sampling errors, etc. Therefore, to ensure that VaR models are reliable, it is crucial to perform rigorous backtesting. Backtesting is the process of comparing the VaR predictions against actual observed outcomes to determine the accuracy of the model. Since a VaR value is fundamentally unobservable, we need to judge the performance of a model by checking whether these estimates are consistent with the realized returns or not, given the confidence level at which these estimates were obtained. Thus, VaR backtesting is critical for evaluating the accuracy of the VaR forecasts.
@@
An alternative to AE ratio is the proportion of failure test (POF), also known as the unconditional coverage test (UC). which statistically evaluates whether the actual number of VaR violations matches the expected number of violations over time, based on the chosen confidence level. For instance, it calculates the failure rate defined as the proportion of the number of violations ($A$) to the total number of observations $T$ and examines its consistency with the model’s confidence level $\alpha$. More specifically, let $I_{t}$ denotes the indicator function that the realized return was less than the VaR estimate, i.e.,
@@
As stated in Kupiec (1995), the UC test is capable of rejecting a model for both low and high failures, but has poor power in general. Further, the UC test tests the number of violations, while it is crucial to test whether these losses are spread evenly over time or not. The Conditional Coverage (CC) test, proposed by Christoffersen (1998), does the same, that is, it tests whether the violations are randomly distributed over time and not clustered. That is, CC test takes into account any conditionality in VaR forecasts: the model must respond to clustering events such as if volatilities are high in some periods and low in others. The test statistics is given by
@@
In the evaluation of VaR forecasts, the quantile score is a commonly used metric and is now standard in the VaR literature (Ehm et al., 2016; Gneiting, 2011; Taylor, 2019; Bams et al., 2017). The above mentioned statistical tests are good to evaluate the statistical adequacy of VaR forecasts, that is, a model for VaR forecasting is deemed to be an adequate model if we fail to reject the null hypothesis. However, these tests cannot conclude whether an ‘adequate’ model outperforms another ‘adequate’ model. Scoring functions allow us to do an honest assessment of VaR as a perfect VaR estimate would always minimize the scoring function. The quantile score $QS_{t}$ can be defined as follows:
@@
For every instant $t$, a model’s score increases either by (a) $\alpha-1$ fraction of the distance between the realized return and the VaR forecast, if realized return is smaller than the VaR forecast for the same period, or (b) by $\alpha$ fraction of the distance otherwise. This scoring rule is consistent for quantile estimation, meaning that it achieves its minimum value when the quantile forecast $q_{t}$ correctly estimates the true quantile of the distribution. Therefore, models that produce lower average quantile scores in the out-of-sample period are considered superior. Finally, to compare the performance of two models, we use the Diebold–Mariano (DM) test (Diebold and Mariano, 2002) with